AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 408     Entry time: Sat Jun 24 10:38:42 2023
Author: JM, TD 
Subject: Section 9 closed, begin pumping  

23/06/23

The CARME chamber and interaction chamber were connected and section 9 has been fully closed. New valves have been installed on the interaction chamber, in addition to new diamond windows. A leak cart has been connected to the backing line of CARME. Before begining the pumpdown, the manual valve between the interaction chamber and the electron target was closed. 

The pumpdown was started at 17:15,  the valve on the scroll pump was very slowly opened to not damage the berylium window. By 18:00 a pressure of 1 mbar was achieved and the turbopump was activated. By 18:20 a pressure of 2E-3 mbar was achieved, the ion gauge was activated and briefly flashed 5E-4 mbar but this is over its limit so the gauge shutdown. At 19:20, 1.3E-3 mbar was achieved, the ion gauge was activated but the reading was 1.7E-4 mbar which is still over its limit. The turbo fan was not working and temperatures on the turbo were ~50C and increasing, so it was decided to close the main gate valve and turno off the turbo overnight in order to fix the fan.

A bias test of the detectors was also performed while in vacuum, T=20.5 C. All detectors show expected V-I curve and the leakage current is ~1-2 uA lower than the previous test in air, probably due to reduced light leaks. The top right detector did initially produce >10 uA at 20 V but the adaptor card was not properly seated on the p-n feed-through. Seating the adaptor card correctly yielded the expected V-I curve. Further inspection of the pins on this feed-through will be performed tomorrow.  

 

24/06/23

Overnight the pressure had increased to 0.8 mbar. Opening the gate valve decreased the pressure by an order of magnitude. The turbo fan was fixed and the turbo was re-activated, the pressure dropped to 1E-3 mbar level within 20 minutes.  Ion gauge was turned on after 30 minutes with a pressure of 3.7e-5 mbar. Pump down plot and comparison to previous pump downs will be uploaded in a seperate elog. Turbo pump maximum temperature reads 46C and has been consistent for several hours. 

V-I tests completed for the right hand side detectors. Leakage currents ~0.4 uA higher than yesterday which is consistent with the increase in temperature of ~ 0.5-0.7C. 

Attachment 1: 23-06-23-VI.png  76 kB  Uploaded Sat Jun 24 17:09:32 2023  | Hide | Hide all
23-06-23-VI.png
Attachment 2: 24-06-23-VI.png  60 kB  Uploaded Sat Jun 24 17:11:09 2023  | Hide | Hide all
24-06-23-VI.png
ELOG V3.1.4-unknown